OpenStudy (anonymous):

11.A cylinder has a radius of 4x + 2 and a height of 5x + 4. Which polynomial in standard form best describes the total volume of the cylinder? Use the formula V = r2h for the volume of a cylinder. 80x3 + 144x2 + 84x + 16 80x3 – 144x2 + 84x + 16 80x3 + 144x2 + 84x – 16 80x3 + 144x2 – 84x + 16 12.A sphere has a radius of 4x + 4. Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S = 4r2 for the surface area of a sphere. 64x2 + 128x + 64 64x2 – 128x + 64 64x2 + 128x – 64 64x2 – 128x – 64

OpenStudy (campbell_st):

so you will have \[V = (4x + 2)^2 \times (5x + 4)\] all you need to do is distribute and collect like terms... but a hint is that very term will be positive... as you will be multiplying postive terms by positive terms... so this may save a little work.

OpenStudy (anonymous):

thx but i still have no clue

OpenStudy (anonymous):

80x3 + 144x2 + 84x + 16 could this possibly be it?

OpenStudy (campbell_st):

ok... so in question 11... in which equation is every term positive...

OpenStudy (campbell_st):

thats it.... every term is positive.... and it saves a bit of work... as you would have had \[(16x^2 + 16x + 4)\times(5x + 4)\]

OpenStudy (campbell_st):

question 12 you would have \[S = 4 \times (4x+4)^2\] all the terms in r are positive... so when you square positives, you get positives... so look for the equation where every term is positive.

OpenStudy (anonymous):

64x2 + 128x + 64?